OpenStudy (anonymous):

Ok so, Find the coordinates of the points of intersection. X-3=0 X^2 + Y^2 = 36 No idea what to do but they tell me the answer is 3, 3 root 3. 0.0? Please help.

OpenStudy (anonymous):

The first equation is a vertical line at x = 3. The second equation is a circle centered at the origin with a radius of 6.

OpenStudy (id21):

x-3=0 x=3 Put it in 2nd equation: \[3^2+y^2=36\] \[y^2=27\] \[y=\pm \sqrt{27}=\pm3\sqrt3\]

OpenStudy (anonymous):

(3, 3 root 3) (3, -(3 root 3))